OpenStudy (anonymous):

if it takes 7 seconds to reach the ground, then its speed at impact is __m/s. the total distance fallen is __m, and its acceleration of fall just before impact is __m/s.

OpenStudy (anonymous):

The two equations you use is v=gt and d = 1/2gt2

OpenStudy (stormfire1):

(excluding drag) For speed at impact: \[V{f}=V{i}+at\]\[Vf = -9.8m/s^2*7s = -68.6 m/s\] For distance: \[Yf = Yi + \frac{1}{2}g*t\]\[Yf = 0m - \frac{9.8m/s^2}{2}(7s)=-34.3m\]It's acceleration is constant throughout so it's just g: \[-9.8m/s^2\]

OpenStudy (stormfire1):

I left out the fact that Vi = 0m/s in the first part...since it's assumed something was dropped from 0 velocity

OpenStudy (stormfire1):

If it actually started out with a velocity, you could just plug it in there.

OpenStudy (anonymous):

So would the first blank be 70m/s, 350m, and 7m/s2

OpenStudy (anonymous):

is that correct

OpenStudy (stormfire1):

Without rounding the blanks would be (in order) -68.6m/s, -34.3m and -9.8m/s^2...they're all negative based on me choosing the y reference being 0 from where the object started to fall. Hopefully that makes sense